# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ I Introduction
Notations: Let $\mathcal{H}$ denote a Hilbert’s space. Let $\mathcal{L}(\mathcal{H})$, $\mathcal{R}(\mathcal{H})$, $\mathcal{P}(\mathcal{H})$ and $\mathcal{D}(\mathcal{H})$ respectively denote the set of all linear, hermitian, positive semidefinite and density operator over the Hilbert’s space $\mathcal{H}$. $\mathrm{I}\in\mathcal{P}(\mathcal{H})\subset\mathcal{R}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ is identity transformation. $\nu_{1}-\nu_{2}\in\mathcal{P}(\mathcal{H})$ is also written as $\nu_{1}\geq\nu_{2}$ or $\nu_{2}\leq\nu_{1}$. Any operator $\Pi\in\mathcal{P}(\mathcal{H})$ is a projection operator if $\Pi\Pi=\Pi$. For any $\nu\in\mathcal{R}(\mathcal{H})$, $\Pi_{\nu}$ denotes projection operator onto the eigenspace spanned by set of eigenvectors corresponding to all non-zero eigenvalues of $\nu$. Note that by the properties of projector $\Pi_{\nu}\nu\Pi_{\nu}=\nu$. $\Pi_{\nu}^{\max}$ and $\Pi_{\nu}^{\min}$ denote the projection operator onto eigen-space corresponding to maximum and smallest non-zero eigenvalues of $\nu$, denoted as $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$ respectively. For some projection operator $\Pi\in\mathcal{P}(\mathcal{H})$, $\mathcal{P}(\Pi)\subset\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\Pi)\subset\mathcal{D}(\mathcal{H})$ denote set of all positive operators and density operators respectively that are invariant w.r.t. $\Pi$ i.e. $\mathcal{P}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{P}(\mathcal{H})\}$ and $\mathcal{S}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{D}(\mathcal{H})\}$. For any $\nu\in\mathcal{P}(\mathcal{H})$, $\nu^{-1}$ is an operator obtained by substituting all non-zero eigenvalues of $\nu$ by their inverse. Note that, with this definition of generalized inverse, we get $\nu^{-1}\nu=\Pi_{\nu}$. Given a pair of operators $\nu_{1},\nu_{2}\in\mathcal{P}(\mathcal{H})$, we denote $R_{\max}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty}$ and $R_{\min}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty,0}$. This is related to the max relative entropy of the states, which is defined as log of maximum eigenvalue of $\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}$ [26].
The paper is organized as follows. Section II describes the system model, the problem statement of testing the two states $\rho$ and $\sigma$, and a derivation of the minimum possible postselected error from the literature before defining the set of error-minimizing measurements and the maximum acceptance over this set. Section III and Section IV contain result for postselected symmetric hypothesis testing problem for the case when $\Pi_{\rho}=\Pi_{\sigma}$ and $\Pi_{\rho}\neq\Pi_{\sigma}$ respectively. In Section III, we begin with deriving the condition on a measurement to achieve the minimum postselected error. Then we characterize the exact set of measurements that achieve minimum postselected error and provide construction for an arbitrary measurement from this set. For better exposition, we give an example to show how the value of acceptance varies over this set and another example to illustrate finding error minimizing measurements. Then, we have maximized acceptance and presented a measurement that achieves the maximizes acceptance. Section IV follow a similar pattern. To maintain the flow, we have just given the key results in the main text while supporting mathematical results are stated and derived in the appendix. Now, we proceed towards describing the system model.
@@ II System Model
We consider a quantum object which can be in one of the two possible states $\rho,\sigma\in\mathcal{D}(\mathcal{H})$. The objective is to determine its state which is unknown. The quantum object being in the state $\rho$ corresponds to null hypothesis and the quantum object being in the state $\sigma$ corresponds to the alternative hypothesis. A measurement is performed with positive operator-valued measure (POVM) $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-(\Lambda_{\rho}+\Lambda_{\sigma})\}$. Note that $\Lambda_{\rho},\Lambda_{\sigma}$ must satisfy the condition $\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})$ and $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ for $\Lambda$ to be a valid measurement. Hence, the set of all such measurements is given by
=$\mathcal{M}\stackrel{{\scriptstyle\Delta}}{{=}}\left\{\Lambda:\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\},\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (1)
On measurement, one of the three outcomes corresponding to the operators $\Lambda_{\rho}$, $\Lambda_{\sigma}$ or $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$ is obtained. The first outcome (corresponding to the operator $\Lambda_{\rho}$) corresponds to accepting the null hypothesis and declaring the unknown state as $\rho$, the second outcome (corresponding to the operator $\Lambda_{\sigma}$) corresponds to accepting the alternative hypothesis and declaring as $\sigma$. If the third outcome (corresponding to the operator $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$) occurs, both the hypotheses are rejected. The probabilities of the outcomes corresponding to the object being in the state $\rho$ and $\sigma$ is summarized in the table below.
Unknown state Probability of…
declaring $\rho$ declaring $\sigma$ rejection accepting null or alternative
@@
If the unknown state is $\rho$ and declared state is $\sigma$, the error is type-1 error and if the unknown state is $\sigma$ and declared state is $\rho$, the error is type- 2 error as mentioned in the table. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$ (collectively denoted as $p$) such that $p_{\rho}>0,p_{\sigma}>0,p_{\rho}+p_{\sigma}=1$, the total error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$.
We define acceptance as the probability of accepting alternative or null hypothesis, i.e. probability of not getting the third outcome. Hence, the acceptance for the state $\rho$ is
@@
The definition of acceptance in (2) and (3) gives a feeling that the higher the acceptance, the better the measurement is, as it corresponds to the probability of accepting at least one hypothesis. But, observe that the expression of acceptance consists of two terms: the probability of correct estimation and the error. Intuitively, we desire the probability of correct estimation to be as big as possible and the error as low as possible, hence a trade-off. Hence, our objective is to maximize acceptance only after minimizing the error, i.e., to find the measurement having the highest acceptance over all the measurements that have minimum error.
The postselected probability of an event is defined as the probability of the event conditioned on the event that alternative or null hypothesis is accepted. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$, the error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$ and the probability that the null or alternative hypothesis is accepted is $p_{\rho}A_{\rho}(\Lambda)+p_{\sigma}A_{\sigma}(\Lambda)$. So, the postselected symmetric error $e(\Lambda)$ is defined as
@@
Further $\Lambda\in\mathcal{M}$ is called as a measurement or POVM and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ even if it is not mentioned. Similarly, $\Lambda\in\mathcal{O}$ is called as an operator and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma}\}$. We now define the minimum postselected symmetric error.
Definition 1 (Minimum postselected symmetric error).
Given prior probability $p_{\rho}$ and $p_{\sigma}$, minimum postselected symmetric error is defined as the minimum achievable postselected symmetric error probability over all measurements i.e.,
@@
Recall that $\mathcal{M}$ is the set of all measurements. It is simplified as[25, Theorem 6]
@@
Now, from [25, Equation (66)], we know that
$\sup_{\Lambda\in\mathcal{O}}{\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}}=\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right).$ (7)
Here, the function $\Xi\left(\nu_{1},\nu_{2}\right)$ is known as Thompson metric[27] and is defined as
@@
Although the minimum error is given in [25], it is not known how to design measurements to achieve this. In this work, one of the key goals is to characterize an arbitrary measurements, which if performed, obtains the minimum error given by (9). The set of all such error-minimizing measurements is formally defined below.
Figure 2: An illustration showing various sets of measurements. $\mathcal{M}$ is set of all measurements. $\mathcal{E}_{s}(\rho,\sigma,p)$ is set of all error-minimizing measurements. Maximum acceptance measurements comprise the set that achieve maximum acceptance over $\mathcal{E}_{s}(\rho,\sigma,p)$.
Definition 2 (Error-minimizing measurements).
The set of all error minimizing measurements is defined as
@@
Within the set of error-minimizing measurements, there are measurements where acceptance for the state $\rho$ and $\sigma$ is maximized (see Fig. 2). The maximum acceptance, denoted as $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, is defined below.
Definition 3 (Maximum acceptance).
@@
We further notice that $e_{s}(\rho,\sigma,p)$ vanishes for the states $\rho$, $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$. So, we split the analysis into two parts discussed in next two sections. The first part considers the case $\Pi_{\rho}=\Pi_{\sigma}$ and second part covers the remaining cases. The next section describes the results for the first part.
III The possible states $\rho$ and $\sigma$ have the same support i.e. $\Pi_{\rho}=\Pi_{\sigma}$
We begin the first subsection by finding the condition on the measurement operators that must be satisfied for the minimum error to be achieved. Based on these conditions, we present some key novel properties of the error-minimizing measurements. Further, we write the set of error-minimization measurements in a parameterized form and give a method to construct an arbitrary error-minimizing measurement. Then we give an example in the next subsection to show that the acceptance for an arbitrary measurement varies with the parameters, although all of them being the error-minimizing operators, thus showing the need to maximizing the acceptance. In the last subsection, we give the expression for maximum achievable acceptance for an arbitrary error-minimizing measurement. The following subsection begins with the characterization of error-minimizing measurements.
III-A The set of all error minimizing measurements and an arbitrary construction
The following theorem derives the condition on measurement operators to achieve the minimum postselected symmetric error, along with providing a novel proof of minimum postselected symmetric error.
Theorem III.1.

### v2

@@ I Introduction
Notation: Let $\mathcal{H}$ denote a Hilbert’s space. Let $\mathcal{L}(\mathcal{H})$, $\mathcal{R}(\mathcal{H})$, $\mathcal{P}(\mathcal{H})$ and $\mathcal{D}(\mathcal{H})$ respectively denote the set of all linear, Hermitian, positive semidefinite and density operator over Hilbert’s space $\mathcal{H}$. $\mathrm{I}\in\mathcal{P}(\mathcal{H})\subset\mathcal{R}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ is identity transformation. $\nu_{1}-\nu_{2}\in\mathcal{P}(\mathcal{H})$ is also written as $\nu_{1}\geq\nu_{2}$ or $\nu_{2}\leq\nu_{1}$. Any operator $\Pi\in\mathcal{P}(\mathcal{H})$ is a projection operator if $\Pi\Pi=\Pi$. For any $\nu\in\mathcal{R}(\mathcal{H})$, $\Pi_{\nu}$ denotes projection operator onto the eigenspace spanned by the set of eigenvectors corresponding to all non-zero eigenvalues of $\nu$. Note that by the properties of projector $\Pi_{\nu}\nu\Pi_{\nu}=\nu$. $\Pi_{\nu}^{\max}$ and $\Pi_{\nu}^{\min}$ denote the projection operator onto eigen-space corresponding to the maximum and smallest non-zero eigenvalues of $\nu$, denoted as $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$ respectively. For some projection operator $\Pi\in\mathcal{P}(\mathcal{H})$, $\mathcal{P}(\Pi)\subset\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\Pi)\subset\mathcal{D}(\mathcal{H})$ denote set of all positive semidefinite operators and density operators respectively that are invariant w.r.t. $\Pi$ i.e. $\mathcal{P}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{P}(\mathcal{H})\}$ and $\mathcal{S}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{D}(\mathcal{H})\}$. For any $\nu\in\mathcal{P}(\mathcal{H})$, $\nu^{-1}$ is an operator obtained by substituting all non-zero eigenvalues of $\nu$ by their inverse. Note that, with this definition of inverse, we get $\nu^{-1}\nu=\Pi_{\nu}$. Given a pair of operators $\nu_{1},\nu_{2}\in\mathcal{P}(\mathcal{H})$, we denote $R_{\max}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty}$ and $R_{\min}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty,0}$. $R_{\max}$ is related to the max relative entropy of the states, defined as the log of the maximum eigenvalue of $\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}$ [11].
Organization: The paper is organized as follows. Section II describes the system model, the problem statement describing PSD between the two states $\rho$ and $\sigma$ as a hypothesis problem, and the minimum possible postselected error from the literature, followed by definitions of acceptance, the set of error-minimizing measurements and the maximum acceptance achieved over this set.
Theorem 1: presents the condition for a measurement to be an error minimizing measurement Theorem 2: gives the set of all error minimizing measurements $\mathcal{E}_{s}$ When the supports of $\rho$ and $\sigma$ are equal i.e. $\Pi_{\rho}=\Pi_{\sigma}$ Theorem 3: evaluates the acceptance for each element in the set $\mathcal{E}_{s}$ Theorem 4: gives the maximizes acceptance achievable by measurement in the set of all error minimizing measurements $\mathcal{E}_{s}$ Section III Theorem 5 : gives the set of all error minimizing measurements $\mathcal{E}_{s}$ When the supports of $\rho$ and $\sigma$ are not equal i.e. $\Pi_{\rho}\neq\Pi_{\sigma}$ Lemma 1,2,3: parameterize an arbitrary error minimizing measurement Theorem 6: evaluates the acceptance for each element in the set $\mathcal{E}_{s}$ Theorem 7: gives the maximizes acceptance achievable by measurement in the set of all error minimizing measurements $\mathcal{E}_{s}$ Section IV
Figure 2: Figure showing connection among theorems in Section III and Section IV
Section III studies the scenario when the supports of the two states $\rho$ and $\sigma$ are equal, i.e., $\Pi_{\rho}=\Pi_{\sigma}$ and presents Theorem 1, 2, 3, and 4. A flow-chart illustrating the contributions of these theorems is included in Fig. 2.
  1. Theorem 1
    states the necessary and sufficient condition that any measurement needs to satisfy to achieve the minimum postselected error. We obtain three different conditions in three different cases depending on the prior probability of the unknown state being $\rho$ (or $\sigma$). We state some important observations regarding the error-minimizing measurements in Remark 1, 2, and 3 and Corollary 1-2. Example 1 illustrates how to obtain error-minimizing measurements utilizing Theorem 1 for a pair of states to be discriminated and further shows that acceptance varies over error-minimizing measurements. Example 2(a) shows that one can obtain the most general error-minimizing measurements from the conditions obtained in Theorem 1.
  2. Theorem 2
    presents the set $\mathcal{E}_{s}$ of all error-minimizing measurements in a form that can be parameterized easily for all three cases separately. We derive Theorem 2 by finding the general subspace where the measurements must lie, using the conditions obtained in Theorem 1. This result is a generalization from the literature in the sense that achievability was shown in the literature [33], but an exhausting set of all such measurements was not known. Then, we give a parametric method to construct an arbitrary error-minimizing measurement by varying certain parameters. The method for all the three cases is summarized in Table III. Example 2(b) illustrates illustrates how to obtain error-minimizing measurements for a pair of quantum states and develops an intuition towards the presented approach to derive the maximum value of acceptance by maximizing one of the parameters.
  3. Theorem 3
    obtains acceptance for an arbitrary error minimizing measurement from the set $\mathcal{E}_{s}$ which was earlier derived in Theorem 2.
  4. Theorem 4
    derives the maximum value of acceptance over the set of all error-minimizing measurements $\mathcal{E}_{s}$. It provides a closed-form expression of the maximum acceptance. In one case, we have given it as an optimization problem. Remark 8 provides an example of a measurement in $\mathcal{E}_{s}$ that achieves the maximum acceptance. This part of the work as presented in Section III is entirely different than USD and is unrelated to the prior literature.
    Section IV studies USD as a special case of PSD when the supports of the two states $\rho$ and $\sigma$ are not equal, i.e., $\Pi_{\rho}\neq\Pi_{\sigma}$.
  5. Theorem 5
    obtains the set of all error minimizing measurements $\mathcal{E}_{s}$, in a form that is easy to parameterize. We observe that the error may vanish in three different ways, resulting in three different conditions. We express $\mathcal{E}_{s}$ as a union of three sets, each satisfying one of the three conditions. We discuss the properties of the measurements in each set. We then show that two sets remain empty when $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$. Further, these measurements are parameterized in Lemma 1, 2, and 3.
  6. Theorem 6
    obtains the acceptance for an arbitrary error minimizing measurement from the set $\mathcal{E}_{s}$.
  7. Theorem 7
    obtains the maximum acceptance over the set of all error-minimizing measurements $\mathcal{E}_{s}$ by maximizing the acceptance obtained on the three sets, and the result is summarized in Table V in the theorem.
Section V concludes the paper with a summary and possible generalizations of this work.
To maintain the flow, we have just given the key results in the main text while supporting mathematical results are stated and derived in the appendix. Now, we proceed to describe the system model.
@@ II System Model
We consider a quantum object that can be in one of the two possible states $\rho$, or $\sigma\in\mathcal{D}(\mathcal{H})$, with prior probabilities $p_{\rho}$ and $p_{\sigma}$, respectively. The objective of PSD is to determine its state with a high accuracy via an appropriately designed measurement. PSD can be seen as a post-selection quantum hypothesis testing problem where the quantum object being in the state $\rho$ corresponds to the null hypothesis, while the quantum object being in the state $\sigma$ corresponds to the alternative hypothesis. A measurement is performed with positive operator-valued measure (POVM) $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-(\Lambda_{\rho}+\Lambda_{\sigma})\}$. Note that $\Lambda_{\rho},\Lambda_{\sigma}$ must satisfy the condition $\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})$ and $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ for $\Lambda$ to be a valid measurement. Hence, the set of all such measurements is given by
$\mathcal{M}\stackrel{{\scriptstyle\Delta}}{{=}}\left\{\Lambda:\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\},\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (1)
The performed measurement results in one of the three outcomes corresponding to the operators $\Lambda_{\rho}$, $\Lambda_{\sigma}$, or $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$. The first outcome (corresponding to the operator $\Lambda_{\rho}$) corresponds to accepting the null hypothesis equivalent to declaring the unknown state as $\rho$, the second outcome (corresponding to the operator $\Lambda_{\sigma}$) corresponds to accepting the alternative hypothesis and declaring it as $\sigma$. If the third outcome (corresponding to the operator $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$) occurs, no decision is made, rejecting the test as inconclusive. The first two outcomes collectively represent the selection or acceptance of the test, where a decision is taken to select one of the hypothesis. Various outcomes and their probabilities are summarized in Table II.
Actual state of Probability of…
the quantum object declaring $\rho$ declaring $\sigma$ rejection accepting null or alternative
@@
TABLE II: Various outcomes and their probabilities for measurement $\Lambda$ given object is in the state $\rho$ or $\sigma$.
Type-1 error refers to the error when the state of the quantum object is declared to be $\sigma$ when the state is actually $\rho$. Similarly, Type-2 error refers to the error when the state of the quantum object is declared to be $\rho$ when the state is actually $\sigma$. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$ (collectively denoted as $p$) such that $p_{\rho}>0,p_{\sigma}>0$, and $p_{\rho}+p_{\sigma}=1$, the total error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$.
We define acceptance as the probability of accepting the test. Note that this corresponds to selecting one of the first two outcomes and excludes the third outcome. Hence, the acceptance for the state $\rho$ is
@@
It may seem from the definition of acceptance given in (2) and (3) that the higher acceptance leads to better measurement, as it corresponds to the probability of accepting at least one hypothesis. However, we can observe that the expression of acceptance consists of two terms: the probability of correct decision and the probability of error. Intuitively, we desire the probability of correct decision to be as large as possible and the error as low as possible, hence a trade-off. Hence, our objective is to maximize acceptance only after minimizing the error, i.e., to find the measurement with the highest acceptance over the set of all the measurements with the minimum error.
The postselected probability of an event is defined as the probability of the event conditioned on the event that the alternative or the null hypothesis is accepted. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$, the error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$ and the probability that the null or alternative hypothesis is accepted is $p_{\rho}A_{\rho}(\Lambda)+p_{\sigma}A_{\sigma}(\Lambda)$. So, the postselected symmetric error $e(\Lambda)$ is defined as [33]
@@
Further $\Lambda\in\mathcal{M}$ is a measurement or POVM and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. $\Lambda\in\mathcal{O}$ is an operator (pair) and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma}\}$ without additional constraints of being a measurement. We now define the minimum postselected symmetric error.
Definition 1 (The minimum postselected symmetric error).
Given prior probability $p_{\rho}$ and $p_{\sigma}$, the minimum postselected symmetric error is defined as the minimum achievable postselected symmetric error probability over all measurements, i.e.,
@@
Here, $\mathcal{M}$ is the set of all measurements. Utilizing (4), $e_{s}(\rho,\sigma,p)$ can be simplified as[33, Theorem 6]
@@
The last step follows from the fact that, for every element $\Lambda_{o}\in\mathcal{O}$, we have a $\Lambda_{m}\in\mathcal{M}$ such that the expression $\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$ takes the same value for both $\Lambda_{o}$ and $\Lambda_{m}$. Also, for every element $\Lambda_{m}\in\mathcal{M}$, we have a $\Lambda_{o}\in\mathcal{O}$ such that $\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$ takes same value for both $\Lambda_{m}$ and $\Lambda_{0}$. We have written for set $\mathcal{O}$ because we get a simple subspace constraint for obtaining the supremum in (6) and (7), as will be clear from Theorem 1.
Now, from [33, Equation (66)], we know that
$\sup_{\Lambda\in\mathcal{O}}{\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}}=\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right),$ (7)
where the function $\Xi\left(\nu_{1},\nu_{2}\right)$ is Thompson metric [38], is given as
@@
Although the minimum error is given in [33], along with an example measurement achieving this error, it is not known how to design an arbitrary error-minimizing measurement. One of the key goals in this work is to characterize the complete set of measurements with postselected error to have the minimum value as given in (9), which we define next.
Figure 3: An illustration showing various sets of measurements. $\mathcal{M}$ is set of all measurements. $\mathcal{E}_{s}(\rho,\sigma,p)$ is set of all error-minimizing measurements. Maximum acceptance measurements comprise the set that achieves the maximum acceptance over $\mathcal{E}_{s}(\rho,\sigma,p)$.
Definition 2 (Set of error-minimizing measurements).
The set of all error-minimizing measurements is defined as
@@
Within the set of error-minimizing measurements, there are measurements for which the acceptance for the state $\rho$ and $\sigma$ is maximized (see Fig. 3). The maximum acceptance is defined below.
Definition 3 (The maximum acceptance).
@@
For the measurements that achieve the minimum error, we are interested in understanding how frequently a decision is made for each possible input state. Hence, we define acceptance of each state individually. The benefit of defining in this fashion is evident from Theorem 6, where acceptance for one of the states turns out to be zero for error-minimizing measurements. This observation would not be evident otherwise.
It can be observed from the expression of the minimum error in (9) and the definition of $\Xi(\cdot,\cdot)$ in (8) that a finite non-zero minimum error exists in the case when the supports of the two states are equal, i.e., $\Pi_{\rho}=\Pi_{\sigma}$. We already know from Theorem 2 in [13] that such cases cannot be discriminated unambiguously. Error minimizing measurements for such cases ($\Pi_{\rho}=\Pi_{\sigma}$) are discussed in Section III. In the remaining cases, the error vanishes, and the analysis reduces to USD. Such cases are discussed in Section IV. The analysis is split this way because the two cases have different subspace constraints on the error-minimizing measurements, thus requiring separate treatments.
III Case: $\rho$ and $\sigma$ have the same support i.e. $\Pi_{\rho}=\Pi_{\sigma}$
We now study PSD for the case where the supports of two states are not equal, and hence, USD is not feasible. We begin with the first subsection by finding the condition on the measurement operators that must be satisfied for the minimum error to be achieved. We start with a key step (7) in the derivation of (9) from [33], building towards a direction for obtaining the error-minimizing condition on any measurement. We note that any measurement that achieves the supremum in (7) is error-minimizing. We focus on obtaining the condition for achieving this supremum in the proof of Theorem 1 given in Appendix A. Although the objective of Theorem 1 is to get the condition for an arbitrary measurement to be an error-minimizing measurement, we also obtained a novel derivation of (9), which is included as a result of the same theorem. Next, we present some key novel properties of the error-minimizing measurements based on the derived condition. Further, we write the set of error-minimization measurements in a parameterized form, providing a method to construct an arbitrary error-minimizing measurement. Then, we provide an example to show that the acceptance for an arbitrary error-minimizing measurement varies with these parameters, even though each one achieves the minimum error. This demonstrates the need to maximize acceptance for obtaining an optimal measurement. In the last subsection, we derive the expression for maximum achievable acceptance over the set of all error-minimizing measurements.
III-A The set of all error minimizing measurements and construction of an arbitrary error-minimizing measurement
Recall that the minimum postselected error in (9) is obtained in terms of $\Xi(\cdot,\cdot)$. From (8), note that, if $\Pi_{\rho}=\Pi_{\sigma}$, we get $\Xi(p_{\rho}\rho,p_{\sigma}\sigma)=\max(R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho),R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho))$. As we see later that the analysis depends on the relative value of the two terms $\max(R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, and $R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho))$, we will study the following three cases separately.
  1. $\bullet$ Case $\underline{\mathcal{C}\mathrm{1}}$:
    $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)>R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$,
  2. $\bullet$ Case $\underline{\mathcal{C}\mathrm{2}}$:
    $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)<R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, and
  3. $\bullet$ Case $\underline{\mathcal{C}\mathrm{3}}$:
    $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$.
In particular, the proof of Theorem 1 shows that we obtain distinct conditions for an error-minimizing measurement for each of these three cases. Note that the value of two possible states and their prior probabilities determine which case the analysis would correspond to. The following theorem derives the condition on the measurement operators to achieve the minimum postselected symmetric error, along with providing a novel proof of deriving its minimum possible value.
Theorem 1.